OpenStudy (anonymous):

What is the simplified form of x plus 1 over x squared plus x minus 6 ÷ x squared plus 5x plus 4 over x plus 4 ? 1 over the quantity x plus 3 times the quantity x plus 4 1 over the quantity x plus 3 times the quantity x minus 2 1 over the quantity x plus 4 times the quantity x minus 2 1 over the quantity x plus 3 times the quantity x plus 1

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\frac{ x + 1 }{ x^2 + x -6 } \div \frac{ x^2 + 5x + 4 }{ x + 4 }\] @satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you go that right?

OpenStudy (anonymous):

nice use of the equation tool btw

OpenStudy (anonymous):

i can show you another trick in using wolfram

OpenStudy (anonymous):

okay and heres the question

OpenStudy (anonymous):

simplify (12z^2 - 25z + 12)/ (3z^2 + 2z - 8)

OpenStudy (anonymous):

using the link i got (4z - 3)/(x + 2)

OpenStudy (anonymous):

ok i so take what you wrote, copy and paste it directly in to wolfram

OpenStudy (anonymous):

yeah i did that and got (4z - 3)/(x + 2)

OpenStudy (anonymous):

yeah i get that too, but of course with a z

OpenStudy (anonymous):

i could also show you how to do it without wolfram, it is "factor and cancel"

OpenStudy (anonymous):

oh yeah typo lol sorry okay i think it'll be better without wolframe alpha in case my teacher asks

OpenStudy (anonymous):

ok then here is the deal do you know how to factor?

OpenStudy (anonymous):

sometimes but not really it'll be helpful if you showed me though

OpenStudy (anonymous):

i don't actually, believe it or not if i wanted to factor i would cheat

OpenStudy (anonymous):

what i mean is that i have no good way to show you how

OpenStudy (anonymous):

i usually use a site for factoring

OpenStudy (anonymous):

that works, of course wolfram will do that too

OpenStudy (anonymous):

so in case the teacher asks, you can say something like this ;

OpenStudy (anonymous):

\[ (12z^2 - 25z + 12)=(4 z-3) (3 z-4)\] and \[(3z^2 + 2z - 8)= (3 z-4) (z+2)\]

OpenStudy (anonymous):

so \[\frac{12z^2+25z+12}{3z^2+2z-8}=\frac{4z-3)(3z-4)}{(3z-4)(z+2)}\]

OpenStudy (anonymous):

oh and then you cross out (3z - 4)?

OpenStudy (anonymous):

there is a common factor top and bottom of \(3z-4\) which cancels, you get \[\frac{(4z-3)(3z-4)}{(3z-4)(z+2)}=\frac{4z-3}{z+2}\]

OpenStudy (anonymous):

exactly what you said

OpenStudy (anonymous):

that is the idea of all of these factor and cancel

OpenStudy (anonymous):

okay cool i get it now

OpenStudy (anonymous):

as to how to factor, you are pretty much on your own, but you can use wolfram to do it for sure

OpenStudy (anonymous):

okay thnx

OpenStudy (anonymous):

and you can use wolfram to give the final answer too, but once you have it factored it is easy to cancel

OpenStudy (anonymous):

ill also ask someone else on os

OpenStudy (anonymous):

want to try another one ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok i get a cup of coffee, you post

OpenStudy (anonymous):

ok What polynomial identity should be used to prove that 21 = 25 - 4?

OpenStudy (anonymous):

wow do you have any choices? \(21=25-4\) is simple arithmetic

OpenStudy (anonymous):

What polynomial identity should be used to prove that 21 = 25 - 4? Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (anonymous):

ooh ok

OpenStudy (anonymous):

25 and 4 are two squares

OpenStudy (anonymous):

then B?

OpenStudy (anonymous):

\[25=5^2\] and \[2^2\] so yeah B

OpenStudy (anonymous):

what a dumb retricequestion next?

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

given the parent function of f(x) = x3, what change will occur when the function is changed to f(x - 3)? Shift to the right 3 units Shift to the left 3 units Shift up 3 units Shift down 3 units

OpenStudy (anonymous):

when you subtract inside the function it move is 3 units to the RIGHT

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

ased on the table of values below, find the slope between points where x = 1 and where x = 4. x y 1 8 3 6 4 −1

OpenStudy (anonymous):

what is the corresponding y value when x = 1?

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

8

OpenStudy (anonymous):

\[\frac{-1-8}{4-1}\] is whatyou have to compuote

OpenStudy (anonymous):

i get \[\frac{-9}{3}=-3\]

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

how many more?

OpenStudy (anonymous):

a couple Which of the following expressions is the conjugate of a complex number with 2 as the real part and 3i as the imaginary part? 2 + 3i 2 − 3i 3i + 2 3i − 2

OpenStudy (anonymous):

the conjugate of \(a+bi\) is \(a-bi\)

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

the conjugate of \(2+3i\) is \(2-3i\) that was an easy one

OpenStudy (anonymous):

yeah b

OpenStudy (anonymous):

okay thnx

OpenStudy (anonymous):

yw next?

OpenStudy (anonymous):

Two car washers, Michelle and Nancy, are working on your car. Michelle can complete the work in 6 hours, while Nancy can complete the work in 3 hours. How many hours does the car washing take if they work together? 3 2 six eighths four thirds

OpenStudy (anonymous):

we go right to the answer (otherwise it take like forever) but remember this simple trick \[\frac{6\times 3}{6+3}\]

OpenStudy (anonymous):

got that ?

OpenStudy (anonymous):

so 18/9 = 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Bob and Susie wash cars for extra money over the summer. Bob's income is determined by f(x) = 6x + 13, where x is the number of hours. Susie's income is g(x) = 4x + 18. If Bob and Susie were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Bob works 3 hours. Create the function h(x) and indicate if Bob will make more money working alone or by teaming with Susie. h(x) = 2x + 5, work alone h(x) = 2x + 5, team with Susie h(x) = 10x + 31, team with Susie h(x) = 10x + 31, work alone

OpenStudy (anonymous):

see the trick? it is easy

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[g(x) = 4x + 18\] \[h(x)=6x+13\] \[g(x)+h(x)=4x+18+6x+13\]

OpenStudy (anonymous):

combine like terms,what do you get?

OpenStudy (anonymous):

10x + 31

OpenStudy (anonymous):

ok good so C or D

OpenStudy (anonymous):

yup but i think its with team work so C?

OpenStudy (anonymous):

not sure if bob works 3 hours alone he makes \[f(3)=6\times 3+13=18+13=31\]

OpenStudy (anonymous):

oh so D

OpenStudy (anonymous):

don't jump the gun

OpenStudy (anonymous):

if they both work 3 hours they make a total of \pg(3)=10\times 3+31=61\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[h(3)=10\times 3+31=61\]\]

OpenStudy (anonymous):

if he has to split it, then yes he is better off working alone

OpenStudy (anonymous):

so is its C then?

OpenStudy (anonymous):

i think D not really clear, but i think D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (3, 6) and a directrix of y = 4?

OpenStudy (anonymous):

this takes a second

OpenStudy (anonymous):

k

OpenStudy (anonymous):

half way between 6 and 4 is 5, so the vertex is \((3,5)\)

OpenStudy (anonymous):

therefore it will look like \[4p(y-5)=(x-3)^2\] we need \(p\)

OpenStudy (anonymous):

the distance between 4 and 5 is 1, so p = 1 and one answer is \[4(y-5)=(x-3)^2\] now sure what your answer choices look like

OpenStudy (anonymous):

f(x) = one fourth (x − 3)2 + 1 f(x) = one fourth (x − 3)2 + 5 f(x) = −one fourth (x − 2)2 + 5 f(x) = −one fourth (x − 2)2

OpenStudy (anonymous):

\[4(y-5)=(x-3)^2\\ y-5=\frac{1}{4}(x-3)^2\\ y=\frac{1}{4}(x-3)^2+5\]

OpenStudy (anonymous):

thnx Which of the following is a solution of x2 + 2x + 8? 4 + i times the square root of 7 −1 + i times the square root of 7 2 + i times the square root of 7 −4 + i times the square root of 7